A certain part of an aircraft engine has a volume of 1.1 ft3. (a) Find the weight of the piece when it is made of lead. (b) If the same piece is made of aluminum, what is its weight? Determine how much weight is saved by using aluminum instead of lead.
Answer #1

First let us convert the volumr of aircraft engine into

since

(a) Density is defined as,

Density = Mass/Volume

Mass = Density*Volume

Density of lead =

Mass of aircraft engine= 11340*0.033 kg

= 374 kg

Weight of engine made of lead = mass*g      ( g is acceleration due to gravity)

= (374*10) N      ( g = )

= 3740 N

(b) Density of aluminium =

Mass of aircraft engine = 2700*0.033 kg

= 89 kg

Weight of engine = 89*10

= 890 N

Weight saved by using aluminium instead of lead = 3740 - 890 N

= 2850 N